# On the absence of anomalous dissipation for the Navier-Stokes equations with Navier boundary conditions: a sufficient condition

### v1

@@ 1. Introduction
(7) $(\omega_{\nu}\times n)_{\tau}+2\lambda(\nu,x)\cdot u_{\nu}-2\big{(}D(n)\cdot u_{\nu}\big{)}_{\tau}=0\quad\hbox{ on }\partial\Omega.$
If one chooses $\lambda=D(n)$ (which is not necessarily nonnegative semidefinite), the Navier boundary condition (5) is equivalent to the following vorticity boundary condition (in three dimensions)
@@
We will utilise the results from [4] in this paper to show a sufficient condition for the convergence of weak solutions of the Navier-Stokes equations to a weak solution of the incompressible Euler equations and the absence of anomalous energy dissipation. Most importantly, we observe that in [4] it was established that the boundary condition for the pressure for weak solutions of the Euler equations, which is rigorously derived in [4] from the weak formulation, is different from the standard classical one for strong solutions. The nonequivalence is further demonstrated by means of an explicitly constructed example.
The literature regarding the dissipation anomaly (and the associated Onsager conjecture) as well as the vanishing viscosity limit is very substantial. Consequently, the following overview will therefore be necessarily incomplete in reviewing all the relevant work. In the case of no-slip Dirichlet boundary conditions for the Navier-Stokes equations, the vanishing of the total (viscous) energy dissipation (as the viscosity goes to zero) in a thin boundary layer (the so-called Prandtl-Von Karman layer) is equivalent to the validity of the vanishing viscosity limit. That is, the solutions of the Navier-Stokes equations converge, as the viscosity tends to zero, to the corresponding solution of the Euler equations, under the assumption of the existence of a strong solution of the Euler equations. This result is known as the Kato criterion and was established in [26].
@@
It is worth stressing that many results have been obtained regarding the vanishing viscosity limit of the Navier-Stokes equations with Navier boundary conditions, see, for example, [2, 13, 10, 23, 27, 3, 32, 35, 34], and references therein. Furthermore, results on the boundary layer expansion for the case of Navier boundary conditions can be found in [40, 25]. Finally, we remark that the study of the vanishing viscosity limit of the Navier-Stokes equations in the case of no-slip boundary conditions has received significant attention in the literature, results in this direction can be found in [37, 29, 16, 33, 5, 31, 14, 38], and references therein. An overview of results for the no-slip case can be found in [34].
@@ 2. Preliminary results and weak formulation of the Euler and Navier-Stokes equations
First we will give a definition of a Leray-Hopf weak solution for the Navier-Stokes equations (1) subject to Navier boundary conditions (4) and (5), which will be based on [28] (the definition for the case of Dirichlet boundary conditions can for example be found in [15]).
@@
    $+\int_{\Omega}u_{\nu,0}\cdot\Psi(x,0)dx-2\int_{0}^{T}\int_{\partial\Omega}(\lambda(\nu,x)\cdot u_{\nu})\cdot\Psi dxdt=0.$
@@
To the knowledge of the authors, the global existence of Leray-Hopf weak solutions for the incompressible Navier-Stokes equations with Navier boundary conditions of the type (4)-(5) has not been established in the literature to this degree of generality. The case $\lambda\equiv 0$ was treated in [22, 41]. In the case $\lambda=\tilde{\lambda}I$ (for some scalar function $\tilde{\lambda}(\nu,x)$), a sketch of the proof of the global existence of Leray-Hopf weak solutions was given in [25]. A full proof of this case was given in [28]. Results in the 2D setting can be found in [13]. We leave the proof of global existence of Leray-Hopf weak solutions for the boundary conditions (4)-(5) to the reader, as it can be established by following the approach in [28] (see also [23]) for instance.
@@ 3. Main result and its proof
Let $\Omega$ be an open set such that $\partial\Omega\in C^{4}$ and $\alpha>0$, and let $\{u_{\nu_{k}}\}$ be a sequence of Leray-Hopf weak solutions of the Navier-Stokes equations (1) (for a sequence of viscosities $\nu_{k}\rightarrow 0$) which obey the Navier boundary conditions (4)-(5) such that the corresponding initial data $u_{\nu_{k},0}$ converge strongly in $C^{0,\alpha}(\Omega)$ to $u_{0}$ (as $\nu_{k}\rightarrow 0$). Moreover, assume that the family of solutions $\{u_{\nu_{k}}\}$ satisfies the following bounds (uniformly as $\nu_{k}\rightarrow 0$)
@@
Moreover, if $\alpha>\frac{1}{3}$ the weak-* limit $u$ above conserves energy so that

### v2

@@ 1. Introduction
(7) $(\omega_{\nu}\times n)_{\tau}+2\lambda(\nu,x)\cdot u_{\nu}-2\big(D(n)\cdot u_{\nu}\big)_{\tau}=0\quad\hbox{ on }\partial\Omega,$
where $\omega_{\nu}=\nabla\times u_{\nu}$ is the vorticity. If one chooses $\lambda=D(n)$ (which is not necessarily nonnegative semidefinite), the Navier boundary condition (5) is equivalent to the following vorticity boundary condition (in three dimensions)
@@
We will utilise the results from [4] in this paper to show a sufficient condition for the convergence of Leray-Hopf weak solutions of the Navier-Stokes equations to a weak solution of the incompressible Euler equations and the absence of anomalous energy dissipation. Most importantly, we observe that in [4] it was established that the boundary condition for the pressure for weak solutions of the Euler equations, which is rigorously derived in [4] from the weak formulation, is different from the standard classical one for strong solutions. The nonequivalence is further demonstrated by means of an explicitly constructed example in [4].
The literature regarding the dissipation anomaly (and the associated Onsager conjecture) as well as the vanishing viscosity limit is very substantial. Consequently, the following overview will therefore be necessarily incomplete in reviewing all the relevant work. In the case of no-slip Dirichlet boundary conditions for the Navier-Stokes equations, the vanishing of the total (viscous) energy dissipation (as the viscosity goes to zero) in a thin boundary layer (the so-called Prandtl-Von Karman layer) is equivalent to the validity of the vanishing viscosity limit. That is, the solutions of the Navier-Stokes equations converge, as the viscosity tends to zero, to the corresponding solution of the Euler equations, under the assumption of the existence of a strong solution of the Euler equations. This result is known as the Kato criterion and was established in [26] (see also the survey [8]).
@@
It is worth stressing that many results have been obtained regarding the vanishing viscosity limit of the Navier-Stokes equations with Navier boundary conditions, see, for example, [2, 13, 10, 23, 27, 3, 32, 35, 34], and references therein. Furthermore, results on the boundary layer expansion for the case of Navier boundary conditions can be found in [40, 25]. Finally, we remark again that the study of the vanishing viscosity limit of the Navier-Stokes equations in the case of no-slip boundary conditions has received significant attention in the literature, results in this direction can be found in [37, 29, 16, 33, 5, 31, 14, 38], and references therein. An overview of results for the no-slip case can be found in [34].
@@ 2. Preliminary results and weak formulation of the Euler and Navier-Stokes equations
First we will give a definition of a Leray-Hopf weak solution for the Navier-Stokes equations (1) subject to Navier boundary conditions (4) and (5), which will be based on [28] (the definition for the case of Dirichlet boundary conditions can for example be found in [15]). In the following definition we take into account the fact that for any divergence-free distributional vector field $v$ one has the following equality in the sense of distributions: $\Delta v=2\nabla\cdot D(v)$.
@@
    $+\int_{\Omega}u_{\nu,0}\cdot\Psi(x,0)dx-2\nu\int_{0}^{T}\int_{\partial\Omega}(\lambda(\nu,x)\cdot u_{\nu})\cdot\Psi dxdt=0.$
@@
To the best knowledge of the authors, the global existence of Leray-Hopf weak solutions for the incompressible Navier-Stokes equations with Navier boundary conditions of the type (4)-(5) has not been established in the literature to this degree of generality. The case $\lambda\equiv 0$ was treated in [22, 41]. In the case $\lambda=\tilde{\lambda}I$ (for some scalar function $\tilde{\lambda}(\nu,x)$), a sketch of the proof of the global existence of Leray-Hopf weak solutions was given in [25]. A full proof of this case was given in [28]. Results in the 2D setting can be found in [13]. For the sake of completeness, in Appendix A we will provide an outline of the proof of the global existence of Leray-Hopf weak solutions for the boundary conditions (4)-(5), by following the approach in [28] (see also [23]).
@@ 3. Main result and its proof
Let $\Omega$ be an open set such that $\partial\Omega\in C^{4}$ and $\alpha\in(0,1)$, and let $\{u_{\nu_{k}}\}$ be a sequence of Leray-Hopf weak solutions of the Navier-Stokes equations (1) (for a sequence of viscosities $\nu_{k}\rightarrow 0$) which obey the Navier boundary conditions (4)-(5) such that the corresponding initial data $u_{\nu_{k},0}$ converge strongly in $C^{0,\alpha}(\Omega)$ to $u_{0}$ (as $\nu_{k}\rightarrow 0$). Moreover, assume that the family of solutions $\{u_{\nu_{k}}\}$ satisfies the following bounds (uniformly as $\nu_{k}\rightarrow 0$)
@@
Moreover, if $\alpha\in\left(\frac{1}{3},1\right)$ the weak-* limit $u$ above conserves energy so that
@@ Acknowledgements
Appendix A Global existence of Leray-Hopf weak solutions for generalised Navier boundary conditions
In this appendix, we will provide a sketch of the proof of the global existence of Leray-Hopf weak solutions for the case of the generalised Navier boundary conditions (4) and (5). The proof proceeds in a similar fashion as in [28]. We first state the result in the following theorem.
Theorem 9.
Let $u_{\nu,0}\in L^{2}(\Omega)$ be a divergence-free vector field such that $(u_{\nu,0}\cdot n)\lvert_{\partial\Omega}=0$, and assume that $\lambda_{ij}\in C^{0}(\partial\Omega)$ for $i,j=1,2,3$. Then there exists a Leray-Hopf weak solution $u$ (in the sense of Definition 1) to the Navier-Stokes equations (1) with the generalised Navier boundary conditions (4) and (5).
Proof.
We will prove the existence of Leray-Hopf weak solutions by employing the Galerkin method. We will denote the Leray projection by $\mathbb{P}$, which we recall is the $L^{2}$-orthogonal projection from $L^{2}(\Omega)$ onto the space $H=\{w\in L^{2}(\Omega)\;\lvert\;\nabla\cdot w=0,\;(w\cdot n)\lvert_{\partial\Omega}=0\}$. In the case of the generalised Navier boundary conditions (4) and (5), we recall that the Stokes operator $A$ is then defined as follows
(31) $A\coloneqq-\mathbb{P}\Delta,$
as a map from $W$ to $W^{\prime}$ (where we recall that the space $W$ and its dual $W^{\prime}$ were defined at the beginning of Section 3). In particular, the Stokes operator is defined on $W$ by duality using the following relation (for $\psi_{1},\psi_{2}\in W$)
(32) $\langle A\psi_{1},\psi_{2}\rangle=(\psi_{1},\psi_{2})_{W}\coloneqq\int_{\Omega}D(\psi_{1}):D(\psi_{2})dx+\int_{\partial\Omega}(\lambda(\nu,x)\cdot\psi_{1})\cdot\psi_{2}dx.$
By means of a generalisation of Proposition 3.2 in [28], we know that the space $H$ has a basis $\{w_{j}\}_{j=1}^{\infty}\subset W$ of eigenfunctions of the Stokes operator, which is orthonormal with respect to the $L^{2}(\Omega)$ inner product. Moreover, the set $\{w_{j}\}_{j=1}^{\infty}$ is also a basis for the space $W$ and is orthogonal with respect to the inner product $(\cdot,\cdot)_{W}$ on $W$, that was defined above.
Now we define $P_{N}$ as the $L^{2}$-orthogonal projection of $H$ onto the space $H_{N}\coloneqq\operatorname{span}\{w_{1},\ldots,w_{N}\}$. We consider the following integral formulation of the order $N$ truncated Galerkin approximation system to equation (13)
(33) $\int_{\Omega}\bigg[\partial_{t}u_{\nu}^{N}\cdot\Psi+\nabla\cdot(u_{\nu}^{N}\otimes u_{\nu}^{N})\cdot\Psi+2\nu D(u_{\nu}^{N}):\nabla\Psi\bigg]dx$
$+2\nu\int_{\partial\Omega}(\lambda(\nu,x)\cdot u_{\nu}^{N})\cdot\Psi dx=0,\quad u_{\nu}^{N}\lvert_{t=0}=P_{N}u_{\nu,0}$
for all $\Psi\in H_{N}$. By the Picard-Lindelöf theorem, there exists a local-in-time solution to the Galerkin system (33). By choosing $\Psi=u_{\nu}^{N}$ and using the fact that $\nabla\cdot u_{\nu}^{N}=0$, we deduce the following energy equality
(34) $\frac{1}{2}\frac{d}{dt}\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}+2\nu\lVert D(u_{\nu}^{N})(\cdot,t)\rVert_{L^{2}}^{2}+2\nu\int_{\partial\Omega}\lambda(\nu,x):(u_{\nu}^{N}\otimes u_{\nu}^{N})dx=0.$
Now, by using the fact that the matrix field $\lambda$ is nonnegative semidefinite, we conclude that the term $2\int_{\partial\Omega}\lambda(\nu,x):(u_{\nu}^{N}\otimes u_{\nu}^{N})dx$ above is nonnegative and hence can be removed, which yields
(35) $\frac{1}{2}\frac{d}{dt}\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}+2\nu\lVert D(u_{\nu}^{N})(\cdot,t)\rVert_{L^{2}}^{2}\leq 0.$
Next, we recall Korn’s inequality (see for example Lemma IV.7.6 in [11]), which is given by
(36) $\lVert\nabla v\rVert_{L^{2}}\leq C\lVert D(v)\rVert_{L^{2}}+C\lVert v\rVert_{L^{2}},$
for all $v\in H^{1}(\Omega)$ and for some positive constant $C$. Therefore, by virtue of estimate (35) and inequality (36) we get that
(37) $\frac{1}{2}\frac{d}{dt}\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}+\frac{2\nu}{C}\lVert\nabla u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}\leq 2\nu\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}.$
We therefore obtain that the Galerkin approximations $u_{\nu}^{N}$ are global-in-time, since (37) implies that
(38) $\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}\leq e^{4\nu t}\lVert u_{\nu}^{N}(\cdot,0)\rVert_{L^{2}}^{2}=e^{4\nu t}\lVert u_{\nu,0}^{N}\rVert_{L^{2}}^{2}.$
Hence the $L^{2}(\Omega)$-norm of $u_{\nu}^{N}$ is uniformly bounded on any finite time interval. Furthermore, from inequality (37) and the above estimate we deduce the following estimate (which is uniform in $N$) for a fixed $T>0$
(39) $\lVert u_{\nu}^{N}(\cdot,T)\rVert_{L^{2}}^{2}+\frac{4\nu}{C}\int_{0}^{T}\lVert\nabla u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}dt\leq\big(1+e^{4\nu T}\big)\lVert u_{\nu,0}\rVert_{L^{2}}^{2}.$
By then proceeding in a completely analogous fashion to [28, p. 7–8], from equation (39) we obtain
$\lVert\partial_{t}u_{\nu}^{N}\rVert_{L^{4/3}((0,T);W^{\prime})}$ $\leq C(\nu\lVert u_{\nu}^{N}\rVert_{L^{2}((0,T);W)}+\lVert u_{\nu}^{N}\rVert_{L^{2}((0,T);W)}^{3/2}\lVert u_{\nu}^{N}\rVert_{L^{\infty}((0,T);H)}^{1/2}$
$+\nu\lVert\lambda\rVert_{C^{0}(\partial\Omega)}\lVert u_{\nu}^{N}\rVert_{L^{2}((0,T);W)})$
$\leq\nu C(1+\lVert\lambda\rVert_{C^{0}(\partial\Omega)})\sqrt{\frac{C}{4\nu}}\big(1+e^{2\nu T}\big)\lVert u_{\nu,0}\rVert_{L^{2}}$
(40) $+C\left(\frac{C}{4\nu}\right)^{3/4}\big(1+e^{4\nu T}\big)\lVert u_{\nu,0}\rVert_{L^{2}}^{2},$
where the constant $C$ depends on $T$. Then by using the Banach-Alaoglu theorem and the Aubin-Lions lemma, we know that there exists a subsequence $\{u_{\nu}^{N}\}_{N=1}^{\infty}$ (which we do not relabel) which converges weakly to $u$ in $L^{2}((0,T);H^{1}(\Omega))$, and weakly-* to $u$ in $L^{\infty}((0,T);L^{2}(\Omega))$ and strongly in $L^{2}((0,T);L^{2}(\Omega))$ as $N\rightarrow\infty$. In particular, we obtain the strong convergence of the sequence $u_{\nu}^{N}(\cdot,t)$ to $u_{\nu}(\cdot,t)$ in $L^{2}(\Omega)$ for almost every $t\in[0,T]$. In addition, we have $u\in C_{w}([0,T];L^{2}(\Omega))$. Moreover, we recall the trace inequality (see Theorem III.2.19 in [11])
(41) $\lVert v\rVert_{L^{2}(\partial\Omega)}^{2}\lesssim\lVert v\rVert_{L^{2}(\Omega)}\lVert v\rVert_{H^{1}(\Omega)},$
for all $v\in H^{1}(\Omega)$. It is then clear that the sequence $\{u_{\nu}^{N}\}_{N=1}^{\infty}$ converges strongly to $u$ in $L^{2}((0,T);L^{2}(\partial\Omega))$. It then follows in a straightforward manner that one can pass to the limit $N\rightarrow\infty$ in the following weak formulation (for divergence-free $\Psi\in\mathcal{D}([0,T)\times\overline{\Omega};\mathbb{R}^{3})$ with $(\Psi\cdot n)\lvert_{\partial\Omega}=0$)
(42) $\int_{0}^{T}\int_{\Omega}\bigg[u_{\nu}^{N}\cdot\partial_{t}P_{N}\Psi+(u_{\nu}^{N}\otimes u_{\nu}^{N}):P_{N}\nabla\Psi-2\nu D(u_{\nu}^{N}):P_{N}\nabla\Psi\bigg]dxdt$
$+\int_{\Omega}u_{\nu,0}^{N}\cdot P_{N}\Psi(x,0)dx-2\nu\int_{0}^{T}\int_{\partial\Omega}(\lambda(\nu,x)\cdot u_{\nu}^{N})\cdot P_{N}\Psi dxdt=0,$
and obtain the validity of equation (13) in the limit $N\rightarrow\infty$. Subsequently, by integrating equation (34) in time, we find
$\frac{1}{2}\lVert u_{\nu}^{N}(\cdot,t)\rVert_{L^{2}}^{2}$ $+2\nu\int_{0}^{t}\lVert D(u_{\nu}^{N})(\cdot,t^{\prime})\rVert_{L^{2}}^{2}dt^{\prime}+2\nu\int_{0}^{t}\int_{\partial\Omega}\lambda(\nu,x):u_{\nu}^{N}\otimes u_{\nu}^{N}dxdt^{\prime}$
(43) $=\frac{1}{2}\lVert u_{\nu,0}^{N}\rVert_{L^{2}}^{2}.$
Then finally, by using the aforementioned convergence results of the sequence $\{u_{\nu}^{N}\}_{N=1}^{\infty}$ (in particular the weak-* convergence in $L^{\infty}((0,T);L^{2}(\Omega))$ and the weak convergence in $L^{2}((0,T);H^{1}(\Omega))$ together with the weak continuity of the limit), by taking the limit $N\rightarrow\infty$ in inequality (43), we obtain the energy inequality (14), by following similar arguments as for the three-dimensional Navier-Stokes equations with Dirichlet boundary conditions (see, e.g., [15]). ∎